# mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt

### v3

@@ IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
Summary of RQ1: We find that LLMs are generally effective in detecting vulnerabilities. The effectiveness can be further enhanced when the type of vulnerability in interest is indicated in the prompt.
@@ IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
The evaluation of explanation quality remains an open problem, as there are no definitive standards for what constitutes a ”correct” or ”complete” explanation. This inherent subjectivity poses significant challenges for effective evaluation. Therefore, the evaluation of explanations must be contextualized within the specific goals of the vulnerability explanation task. By summarizing and generalizing the “helpfulness” of the generated text, we assess explanation quality from three critical perspectives: Accuracy, Clarity, and Actionability.
@@
In this study, we evaluate the explanations generated by the fine-tuned LLMVulExp for vulnerability code from the SeVC and DiverseVul test sets, alongside ground truth annotations generated using GPT-3.5. Given the large scale of the dataset and the high cost of evaluation, we employ a random sampling approach based on a 95% confidence level and a 5% confidence interval [45]. This yields 384 and 224 samples from SeVC and DiverseVul, respectively, ensuring statistical validity and representativeness in our evaluation.
Our evaluation includes two parts: (1) Manual review: Two authors, with extensive research experience in software vulnerabilities and development, independently review the sampled data, scoring them based on predefined criteria. Disagreements are resolved through discussion until a consensus is reached. The Cohen’s Kappa value [46] for inter-reviewer agreement is 0.76, indicating substantial consistency and reliability. (2) Automated review using LLM: GPT-3.5 is used for automated evaluation. A detailed prompt is provided, describing the vulnerability explanation evaluation task and the specific criteria. The LLM is instructed to output scores for the three evaluation metrics, with 1 indicating satisfaction and 0 indicating non-satisfaction.
@@
We present the results of our dual evaluation process in Table VII. The suffix Gen. refers to explanations generated by LLMVulExp, while Ann. refers to annotations from GPT-3.5. All-Pos. represents cases meeting all three evaluation metrics. Key findings include:
(1) Accuracy of Explanations: LLMVulExp achieves high accuracy (e.g., over 90.0% for SeVC-Gen. in both manual and automated reviews), pinpointing vulnerability risks and code locations. This underscores the importance of accurate vulnerability type identification.
(2) Clarity of Explanations: The clarity scores are 81.4% for SeVC and 94.1% for DiverseVul, indicating the model effectively reduces cognitive barriers for developers analyzing vulnerabilities.
(3) Actionability of Explanations: Explanations from LLMVulExp include actionable suggestions for code modifications, achieving 93.4% for SeVC and 80.5% for DiverseVul. This highlights the model’s potential to support practical remediation efforts.
(4) Effectiveness of Annotation Method: Results from SeVC-Ann. and DIV-Ann. validate the effectiveness of our annotation method in generating high-quality explanatory data, reducing subsequent annotation costs.
(5) Potential of LLM in Automated Assessment: LLM-based automated evaluations closely align with manual reviews. For All-Pos., manual vs. automated results are 76.0% vs. 74.0% for SeVC and 80.1% vs. 71.9% for DiverseVul. This demonstrates LLMs’ potential to scale evaluation tasks efficiently.
@@
TABLE VII: Proportion (%) of Vulnerability Explanation Review Results (RQ2).
Metric SeVC-Gen. SeVC-Ann. DIV-Gen. DIV-Ann.
@@
Accuracy 91.1 93.1 73.3 94.1
Clarity 81.4 81.7 94.1 98.6
Construct. 93.4 94.6 80.5 83.2
All-Pos. 76.0 76.0 59.7 80.1
@@
Accuracy 90.4 97.6 96.4 96.8
Clarity 74.3 77.2 95.0 96.8
Construct. 83.5 88.6 67.9 71.9
All-Pos. 72.8 74.0 67.9 71.9
1int i; 2for(i = 0; i < 1; i++) 3twoIntsStruct * dataBuffer = (twoIntsStruct *) ALLOCA(100 * sizeof(twoIntsStruct)); 4size_t i; 5for (i = 0; i < 100; i++) 6 dataBuffer[i].intOne = 1; 7 dataBuffer[i].intTwo = 1; 8data = dataBuffer; 9printStructLine(&data[0]); 10 11void printStructLine(const twoIntsStruct * structTwoIntsStruct) { 12 printf($d\n$", structTwoIntsStruct->intOne, structTwoIntsStruct->intTwo); 13} 14delete [] data;
{mdframed}[style=mystyle] [type] pointer
[location] The line ”delete [] data” has a pointer-related issue.
@@
(Analysis:) The issue in this code is related to the ”data” pointer, which is being improperly freed using the ”delete” operator. The ”data” pointer is assigned the address of a memory block allocated on the stack using the ”ALLOCA” function, which allocates memory on the stack instead of the heap. The ”delete” operator is used to free memory allocated on the heap using the ”new” operator, and using it to free memory allocated on the stack may cause undefined behavior or crashes.
@@
(Suggestion:) To fix this issue, the program should use the ”free” function to free memory allocated on the stack, or use the ”new” operator to allocate memory on the heap instead of the ”ALLOCA” function. Additionally, the program should ensure that the ”data” pointer points to a valid memory location before it is used or accessed, and should add proper error handling and validation to ensure that the ”data” pointer behaves correctly and safely in all cases.
TABLE VIII: Example of Generated Explanation on SeVC.
IV-B3 Illustrative Example.
To highlight the importance of evaluating vulnerability explanations, we analyze an example from SeVC. The sample pertains to a pointer-type vulnerability caused by incorrect memory deallocation in C code. As shown in Figure IV-B2, the explanation correctly identifies the issue: the delete operator is used to free stack-allocated memory created by the ALLOCA function, which should instead use the free function. The explanation provides actionable suggestions, such as ensuring pointer safety by using the appropriate deallocation methods. It is clear, logically structured, and actionable, satisfying all three evaluation criteria.
Despite these strengths, some generated explanations exhibit issues, such as vague locations, verbose descriptions, or lack of actionable suggestions. While rare, hallucinations may occur, typically involving memorized examples or misinterpreted contexts. Overall, the fine-tuned model demonstrates robust capabilities for generating high-quality vulnerability explanations.
Summary of RQ2: We find that LLMVulExp is capable of generating vulnerability explanations with high levels of Accuracy, Clarity, and Actionability. Additionally, we explore the potential of leveraging LLMs for automated data annotation, significantly reducing the reliance on manual effort.
@@ IV-C RQ3: How do explanations affect the results of vulnerability detection?
Building on the detection capabilities evaluated in RQ1, we consider the detection task as a foundational step in the broader explanation task. This raises an important question: How does fine-tuning for vulnerability explanation influence the detection performance of LLMs compared to models fine-tuned solely for detection? In this RQ, we experimentally examine the impact of incorporating explanatory information on vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory data.
Experimental Setup. We conduct ablation studies under the multi-type vulnerability detection scenario in SeVC (RQ1.2) and the multi-label CWE detection scenario in DiverseVul (RQ1.3). To isolate the effect of explanatory information, we transform the vulnerability explanation fine-tuning task into a detection-only task by removing explanatory data from annotations and adjusting task descriptions and prompts. The resulting LLMs output only vulnerability types, creating a detection-focused model (referred to as Detector). This is compared against the performance of the original explanation-focused model (referred to as Explainer).
Figure 5: Performance of Explainer and Detector.
Experimental Results. The comparative results between Explainer and Detector are illustrated in Figure 5, where ”Mi” and ”Ma” denote Micro and Macro metrics, respectively. Incorporating explanatory data into the fine-tuning process did not significantly reduce detection performance. For instance, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while both achieve 74.8% on DiverseVul. Interestingly, certain metrics improved when explanatory data was included, such as the Macro-F1 score on DiverseVul, where the Explainer outperformed the Detector by nearly 2% (73.2% vs. 71.4%).
@@
Summary of RQ3: Integrating vulnerability explanations into the fine-tuning process preserves detection capabilities and may even enhance performance in specific vulnerability scenarios.
IV-D RQ4: How does the key code extraction impact detection performance?
In this RQ, we investigate whether LLMs can identify key code segments prone to vulnerabilities and enhance detection by focusing on these critical sections. To explore this, we annotate key code segments in the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the annotated key code information.
Experimental Setup. As outlined in Section III, we employ GPT-3.5 to annotate key code segments closely related to vulnerabilities. To avoid label leakage, vulnerability-type tags are concealed during the annotation process. For training samples where key code information was unavailable, the annotations are nullified. Similarly, test samples without key code annotations are excluded from the evaluation. During fine-tuning and inference, the prompts are modified to emphasize task descriptions and instructions, guiding the model to focus on the extracted key code for vulnerability detection.
TABLE IX: F1 of Key Code Extract. on SeVC: (RQ4).
Type Single-Type(RQ1.1) Multi-Type(RQ1.3)

### v4

@@ IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
IV-A5 RQ1.4: Performance on the PrimeVul Dataset
Task Setup. To further probe the limits of current models, we conduct a more stringent evaluation using the PrimeVul dataset. We This task is formulated as multi-class classification, where each vulnerable function belongs to exactly one CWE type. PrimeVul is recognized as a more challenging benchmark due to its wider variety of vulnerability types and more realistic, imbalanced data distribution, which rigorously tests a model’s ability to distinguish between fine-grained vulnerability patterns.
Experimental Results. The performance on PrimeVul, detailed in Table VII, presents a complex and challenging picture that highlights the critical issue of class imbalance in vulnerability detection.
In the multi-class CWE detection task, the high Weighted-F1 scores (e.g., 94.9% for LLMVulExp) initially suggest strong performance. However, these scores are misleading as they are dominated by the models’ accuracy on the major non-vulnerable code. The performance of all models plummets on the Macro-averaged metrics, with the top Macro-F1 score being a mere 21.7% from CausalVul, and our model achieving a comparable 20.7%. This dramatic gap between Weighted and Macro scores proves that all models, while able to learn common vulnerability patterns, largely fail to generalize to the long tail of less frequent CWEs—a critical failure for real-world applications.
This difficulty is further confirmed by the binary detection task. On this more challenging dataset, all models experienced a drastic performance drop compared to their results on DiverseVul. Our model, LLMVulExp, while achieving the highest F1-Score at 39.8%, still operates at a level far below what is required for reliable, practical use.
The findings from PrimeVul strongly suggest that current fine-tuning methodologies for LLMs, despite their promise, are not yet robust enough for complex, imbalanced, real-world vulnerability analysis. This underscores a critical research gap and motivates the need for more advanced techniques to enhance model generalization and robustness against low-frequency data.
TABLE VII: Performance Comparison on PrimeVul (RQ1.4). All metric values are percentages (%). Best results in each row are in bold. * means indicates statistically significant improvement over the baseline method through paired t-test ($p<0.05$).
Task Metric Ours CodeBERT CodeT5 LineVul CausalVul
Multi-Class CWE Detection Precision (Weighted) 96.2 96.0 95.9 95.2 96.1
Recall (Weighted) 93.1 90.2 89.2 93.4 90.4
F1 Score (Weighted) 94.9 93.1 91.9 94.1 93.3
Precision (Macro) 19.2 13.7 18.7 14.4 25.0
Recall (Macro) 25.7 20.1 20.4 19.3 19.2
F1 Score (Macro) 20.7 14.1 15.3 14.0 21.7
Binary Vulnerability Detection Precision 29.5 19.6 18.2 25.8 20.2
Recall 61.1 60.1 61.1 47.8 59.4
F1 Score $\textbf{39.8}^{*}$ 29.6 28.0 33.5 30.1
Summary of RQ1: Our evaluation offers a dual perspective on LLMs for vulnerability detection. It confirms their feasibility and effectiveness on established benchmarks like SeVC and DiverseVul. However, it also reveals their current limitations, as all models struggled with the class imbalance and complexity of the more challenging PrimeVul dataset, highlighting a critical area for future research.
@@ IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
The evaluation of explanation quality remains an open problem, as there are no definitive standards for what constitutes a “correct” or “complete” explanation. This inherent subjectivity poses significant challenges for effective evaluation. Therefore, the evaluation of explanations must be contextualized within the specific goals of the vulnerability explanation task. By summarizing and generalizing the “helpfulness” of the generated text, we assess explanation quality from three critical perspectives: Accuracy, Clarity, and Actionability.
@@
In this study, we evaluate the explanations generated by the fine-tuned LLMVulExp for vulnerability code from the DiverseVul and PrimeVul test sets, alongside ground truth annotations generated using Deepseek-V3. Given the large scale of the dataset and the high cost of evaluation, we employ a random sampling approach based on a 95% confidence level and a 5% confidence interval [54]. This yields 224 and 197 samples from DiverseVul and PrimeVul, respectively, ensuring statistical validity and representativeness in our evaluation.
Our evaluation includes two parts: (1) Manual review: Two authors, with extensive research experience in software vulnerabilities and development, independently review the sampled data, scoring them based on predefined criteria. Disagreements are resolved through discussion until a consensus is reached. The Cohen’s Kappa values [55] for inter-reviewer agreement are 0.78 and 0.81 for two datasets, indicating substantial consistency and reliability. (2) Automated review using LLM: Deepseek-V3 is used for automated evaluation. A detailed prompt is provided, describing the vulnerability explanation evaluation task and the specific criteria. The LLM is instructed to output scores for the three evaluation metrics, with 1 indicating satisfaction and 0 indicating non-satisfaction.
@@
We present the results of our dual-rater (Manual and LLM-Automated) evaluation in Table VIII. The column suffixes denote the data source: -Gen refers to explanations generated by our model (LLMVulExp), while -Ann refers to the higher-quality explanations produced by our annotation method. All-Pos. is a strict metric representing cases that satisfy all three criteria (Accuracy, Clarity, and Actionability). Our analysis yields three key findings:
(1) Generated Explanations Show Value, but Annotated Explanations Are Superior. Our model-generated explanations (-Gen) demonstrate practical value, achieving, for instance, a manual All-Positive score of 51.3% on DiverseVul. However, the results consistently show that the explanations produced via our annotation method (-Ann) are of significantly higher quality across all metrics and datasets. For example, on PrimeVul, the manually-assessed All-Positive score for PRI-Ann (90.3%) is more than double that of PRI-Gen (41.6%). This validates the effectiveness of our annotation pipeline in creating a high-quality dataset for fine-tuning and evaluation. The lower scores on PrimeVul compared to DiverseVul also echo our findings from RQ1, confirming it as a more challenging benchmark for both detection and explanation.
(2) Agreement Varies Significantly Between Objective and Subjective Metrics. The Cohen’s Kappa coefficients, which measure the agreement between manual and LLM raters, reveal a critical insight. For the relatively objective Accuracy metric, which assesses factual correctness (e.g., correct location, valid CWE), the agreement is consistently in the moderate to substantial range (Kappa from 0.52 to 0.65). This indicates that LLMs can reliably replicate human judgment when the evaluation criteria are concrete and verifiable.
In stark contrast, the agreement for the more subjective metrics of Clarity and Actionability is significantly lower and more varied. On DiverseVul, the Kappa values are near zero (0.01 to 0.12), signaling almost no agreement beyond chance. While the agreement is better on PrimeVul, it remains modest. This discrepancy strongly suggests that while LLMs are proficient at automating factual checks, they struggle to consistently align with human assessment of subjective qualities like writing style, comprehensibility, and the perceived usefulness of repair suggestions.
(3) LLMs Are Viable for Automating Concrete, Fact-Based Evaluation. Based on the Kappa analysis, we conclude that LLMs are most suitable for automating the evaluation of specific, clearly-defined aspects of an explanation’s quality. Our comparative analysis of inter-rater agreement reveals a clear and practical boundary for using LLMs in evaluation. Compared to human-human agreement, the human-LLM agreement is relatively lower. Crucially, this significant gap is not uniform. The discrepancy is primarily driven by the LLM’s struggle with nuanced, subjective metrics like Clarity and Actionability. The strong agreement on the ‘Accuracy‘ metric validates the potential of LLM-based automation for reliably checking the factual correctness of generated explanations at scale. This allows human experts to focus their limited time on the more nuanced, subjective aspects that still require human oversight. Our dual-validation strategy, therefore, offers a practical framework: leveraging LLMs for scalable, objective assessment while using targeted human review for subjective quality control, creating an efficient and reliable evaluation pipeline.
@@
TABLE VIII: Proportion (%) of Vulnerability Explanation Review Results and Inter-Rater Agreement (RQ2).
Metric DIV-Gen. DIV-Ann. PRI-Gen. PRI-Ann.
@@
Accuracy 57.4 84.3 43.8 93.0
Clarity 96.7 99.0 71.4 97.8
Action. 82.7 83.2 42.7 90.3
All-Pos. 51.3 71.1 41.6 90.3
@@
Accuracy 62.9 89.3 46.5 94.6
Clarity 90.9 97.5 73.5 99.5
Action. 61.4 85.3 47.0 93.5
All-Pos. 59.9 85.3 46.5 93.0
Agreement (Cohen’s Kappa)
Human vs. Human Agreement (per-metric)
Accuracy 0.92 0.90 0.77 0.75
Clarity 0.76 0.66 0.82 0.85
Actionability 0.60 0.60 0.64 0.65
Human vs. LLM Agreement (per-metric)
Accuracy 0.65 0.52 0.60 0.58
Clarity 0.04 0.01 0.57 0.39
Actionability 0.02 0.12 0.54 0.57
Overall Agreement Comparison
Human vs. Human 0.78 0.81
Human vs. LLM 0.38 0.68
IV-B3 Qualitative Analysis of Vulnerability Explanations
Beyond the numerical scores, a qualitative analysis of the generated explanations is crucial to understand why the models perform as they do. By examining both exemplary and unsatisfactory cases, we can uncover the models’ underlying capabilities and current limitations.
An Exemplary Case. To highlight the importance of evaluating vulnerability explanations, we analyze an example from SeVC. The sample pertains to a pointer-type vulnerability caused by incorrect memory deallocation in C code. As shown in Figure IV-B3, the explanation correctly identifies the issue: the delete operator is used to free stack-allocated memory created by the ALLOCA function, which should instead use the free function. The explanation provides actionable suggestions, such as ensuring pointer safety by using the appropriate deallocation methods. It is clear, logically structured, and actionable, satisfying all three evaluation criteria.
[location] The line “delete [] data” has a pointer-related issue.
@@ IV-B3 Qualitative Analysis of Vulnerability Explanations
(Analysis:) The issue in this code is related to the “data” pointer, which is being improperly freed using the “delete” operator. The “data” pointer is assigned the address of a memory block allocated on the stack using the “ALLOCA” function, which allocates memory on the stack instead of the heap. The “delete” operator is used to free memory allocated on the heap using the “new” operator, and using it to free memory allocated on the stack may cause undefined behavior or crashes.
@@
(Suggestion:) To fix this issue, the program should use the “free” function to free memory allocated on the stack, or use the “new” operator to allocate memory on the heap instead of the “ALLOCA” function. Additionally, the program should ensure that the “data” pointer points to a valid memory location before it is used or accessed, and should add proper error handling and validation to ensure that the “data” pointer behaves correctly and safely in all cases.
TABLE IX: Example of Generated Explanation on SeVC.
Insights from Unsatisfactory Explanations. An analysis of the less satisfactory explanations reveals recurring patterns that not only explain the numerical results in Table VIII but also offer crucial insights for future research.
1. Factual Errors vs. Subjective Deficiencies: Explaining the Kappa Discrepancy. The shortcomings in generated explanations can be broadly categorized into two types, which directly explain the variance in inter-rater agreement (Kappa scores).
Factual Inaccuracies: The first category includes objective, verifiable errors. These are the primary reason for low Accuracy scores and include: (a) Incorrect Localization, where the model identifies the wrong line number; (b) Root Cause Misidentification, where it describes a symptom (e.g., a crash) instead of the underlying flaw (e.g., a use-after-free); and (c) Vulnerability Misclassification. Because these errors are factual, both human and LLM raters can assess them with high consistency, leading to the moderate-to-substantial Kappa agreement (0.52-0.65) observed for the Accuracy metric.
Subjective Deficiencies: The second category involves subjective qualities of the explanation. These include: (a) Generic or Missing Repair Guidance, where suggestions are too vague (e.g., “validate all inputs”); and (b) Poor Clarity, due to excessive verbosity or templated, repetitive phrasing. Assessing these qualities is inherently subjective—what one developer finds clear, another may find verbose. This subjectivity is the primary reason for the low-to-minimal Kappa agreement seen for Clarity and Actionability, especially on the DiverseVul dataset.
2. Shallow Pattern Matching vs. Deep Semantic Reasoning: The Core Limitation. A deeper insight is that many of these failures stem from a single core limitation: the model often relies on shallow pattern matching rather than deep, causal reasoning about the code’s execution flow.
This hypothesis explains several observed phenomena. Incorrect Localization often results from matching superficial tokens (e.g., a function name) without understanding the control flow that activates the vulnerability. The model’s tendency to produce templated or generic explanations suggests it is retrieving and adapting a learned pattern rather than reasoning from first principles. Most importantly, this explains why performance was notably worse on the more complex PrimeVul dataset (Table VIII). The vulnerabilities in PrimeVul likely require a more global understanding of inter-procedural dependencies, which is a known weakness of models that rely on local, pattern-based reasoning. The model excels at explaining common vulnerabilities it has seen frequently during training but struggles when faced with novelty or complexity.
Implications. The qualitative analysis reveals the dual nature of current LLMs for vulnerability explanation. They are powerful pattern recognizers, capable of producing excellent explanations for common vulnerabilities in relatively simple code. However, their reliability diminishes when faced with complex or novel scenarios that require deep semantic understanding. This underscores that while LLMs are invaluable assistive tools, they are not yet infallible replacements for human expertise. The path forward for research is clear: we must develop methods that move beyond pattern matching to instill a more robust, causal reasoning capability in these models.
Summary of RQ2: Our analysis reveals a critical insight for automated explanation evaluation: LLMs show strong agreement with human experts on factual Accuracy but diverge significantly on subjective metrics like Clarity and Actionability. This suggests that the most promising path forward is a hybrid evaluation strategy, where LLMs automate scalable, objective checks, while human oversight remains essential for nuanced, subjective quality assessment.
@@ IV-C RQ3: How do explanations affect the results of vulnerability detection?
Building on the detection capabilities evaluated in RQ1, this section investigates a critical question: Does the act of generating explanations help or hinder the primary task of vulnerability detection? To answer this, we experimentally examine the impact of the explanation task by comparing our standard model, which is fine-tuned to both detect and explain, against a variant fine-tuned solely for detection.
Experimental Setup. We conduct this ablation study on the SeVC, DiverseVul and PrimeVul datasets. To isolate the effect of the explanation task, we create a detection-focused counterpart for our model. This is achieved by removing all explanatory text from the training data and adjusting the instruction prompts so that the model’s only goal is to output the vulnerability label(s). We refer to our standard model as the Explainer and its detection-only counterpart as the Detector. All other hyperparameters and training procedures are kept identical.
Figure 5: Performance comparison of the Explainer (trained to explain and detect) and the Detector (trained only to detect) on the DiverseVul and PrimeVul datasets.
Experimental Results. We first show the comparison of Detector and Explainer on easier SeVC for multi-type vulnerability detection in Table X. Our experimental results reveal that detection-only models outperform interpretable models on the SeVC synthetic dataset. This occurs because synthetic vulnerability patterns in SeVC are highly discernible, and interpretation task training directs the model’s focus toward explanation generation, slightly compromising detection performance. The marginal performance discrepancy indicates strong compatibility between these tasks, demonstrating that interpretation tasks can jointly participate in large model fine-tuning alongside detection objectives.”
TABLE X: SeVC Dataset Experiment Results Comparison
Model Mi-Pre. Mi-Rec. Mi-F1 Ma-Pre. Ma-Rec. Ma-F1
SeVC-Explainer 79.9 80.5 80.1 70.0 72.4 70.8
SeVC-Detector 81.7 81.4 81.2 70.6 75.4 72.0
The comparative results of real-world vulnerability detection, shown in Figure 5, reveal a nuanced and dataset-dependent relationship between explanation and detection performance. The impact of learning to explain is not monolithic; instead, it is dictated by the complexity of the underlying task.
On the simpler benchmark (DiverseVul), the detection-only model (‘Detector‘) shows a marginal performance advantage. For instance, the ‘Detector‘ achieves a slightly higher Macro-F1 score than the ‘Explainer‘ (71.0% vs. 69.9%). This suggests that on less complex tasks, forcing the model to also generate explanations may introduce a minor multi-task learning overhead, slightly diluting its focus on the primary detection objective. On the more challenging benchmark (PrimeVul), this trend dramatically reverses. Here, the ‘Explainer‘ model significantly outperforms the ‘Detector‘ across nearly all metrics. The most notable advantage is seen in the Micro-F1 score (93.2% vs. 88.7%), but the trend also holds for the more challenging Macro-F1 score (20.7% vs. 18.2%). This reversal is a critical finding. It strongly suggests that for complex, real-world vulnerabilities with subtle patterns, the task of generating explanations is not a hindrance. Instead, it acts as a form of implicit regularization or deep supervision. By learning to articulate why a piece of code is vulnerable, the model is compelled to develop a more robust and granular understanding of the vulnerability’s root cause. This deeper semantic comprehension, forged through the explanation task, directly enhances its ability to detect those same complex vulnerabilities, providing a performance benefit that far outweighs any multi-task overhead.
@@
Summary of RQ3: The relationship between explanation and detection is nuanced and depends on task complexity. For simpler vulnerabilities, a detection-only focus is marginally better. For complex, realistic vulnerabilities, however, fine-tuning for explanation significantly boosts detection performance by fostering a deeper, more robust understanding of root causes.
IV-D RQ4: How does the key code extraction impact detection and explanation performance?
In this RQ, we investigate whether explicitly guiding an LLM to focus on critical code segments can enhance its performance. To explore this, we creat datasets for SeVC, DiverseVul and PrimeVul where likely vulnerable code snippets (“key code”) are pre-annotated. We then fine-tune and evaluate our model using this focused information.
Experimental Setup. As outlined in Section III, we employ LLMs to annotate key code segments related to vulnerabilities in our datasets. To prevent data leakage, vulnerability type information is concealed during this annotation process. During the fine-tuning and inference stages for the “With Key Code” model, our prompts are modified to instruct the model to specifically leverage these extracted snippets for its analysis. We then compare this model’s detection performance against the baseline model trained “Without Key Code”. We also briefly assess the impact on explanation quality using the same methodology as in RQ2.
TABLE XI: F1 of Key Code Extract. on SeVC: (RQ4).
Type Single-Type(RQ1.1) Multi-Type(RQ1.2)